Thornton Computer Services, Inc. has been in business for six months. The following are basic operating data for that period: Month Nov. Dec. 330 $6,000 $6,800 $13,000 $21,000 $16,000 $16,500 Oct. 320 Service hours Revenue Operating costs 120 136 260 420 $4,300 $5,300 7,100 $11,200 9,100 $10,600 Required a. What is the average service revenue per hour in each month and the overall average for the six-month period? c. Determine the average contribution margin per hour fixed cost and the variable cost per hour Complete this question by entering your answers in the tabs below. Required A Required B Required C e the bigh-low method to estimate the total monthly fixed cost and the variable cost.per hour.... per hour per month Fixed cost Required A Required c >
Answer #1

answer

a Average service revenue per hour
Formula : Revenue/ Service Hours
July Aug. Sept. Oct. Nov. Dec. Total
Service Hours 120 136 260 420 320 330 1586
Revenue 6000 6800 13000 21000 16000 16500 79300
Average service Revenue per hour 50 50 50 50 50 50 50
Overall Average for the six month period service Revenue per hour 79300/1586 = 50

b

Month Operating Cost Service Hours Variable Cost
Oct high 11200 420
July Low 4300 120
Difference 6900 300 (6900/300)        23

Fixed Cost = Operating Cost in October - (Variable cost per unit * Service   hours in October)

= 11200 -23*420 = 1540

C Average contribution Margin Per Hour
July Aug. Sept. Oct. Nov. Dec. Total
Service Hours 120 136 260 420 320 330 1586
Revenue 6000 6800 13000 21000 16000 16500 79300
Variable Cost 2760 3760 5560 9660 7560 9060 38360
Contribution 3240 3040 7440 11340 8440 7440 40940
Contribution per hour 27.00 22.35 28.62 27.00 26.38 22.55 25.81